Some year, Ben's net salary was 1939 €, while the consumer price index (CPI) was 106.5. One year later, his net salary was 2210 €, while the consumer price index (CPI) was 115.2. Calculate the relative real change in Ben's net salary as percentage.


1
Expert's answer
2022-02-15T17:56:21-0500

The relative nominal change in Ben's net salary as percentage is: (2210 - 1939)/1939×100% = 13.98%.

The inflation rate is:

i = (115.2 - 106.5)/106.5×100% = 8.17%.

So, the relative real change in Ben's net salary as percentage is:

13.98 - 8.17 = 5.81%.